Two cars leave towns 760 kilometers apart at the same time and travel toward each other. One car's rate is 12 kilometers per hour more than the other's. If they meet in 4 hours, what is the rate of the faster car?

Answer by rwm(914)

r*t=d
r*4+(r+12)*4=760
4*(2r+12)=760
4*2*(r+6)=760
r+6=760/8
r+6=95
r=89
r+12=101